OpenStudy (jokersmart):

WILL FAN AND MEDAL!!! Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 36 when x = 3. A. y = x2 + 25 B. y = 4x2 C. y =1/4x2 D. y = 12x

OpenStudy (anonymous):

what you think it's is

OpenStudy (jokersmart):

B?

OpenStudy (dtan5457):

Your direct proportion would be. y=k(x^2) You need to solve for k first. 36=(3^2)k k=4

OpenStudy (dtan5457):

Yes, you are right.

OpenStudy (jokersmart):

Awesome! :'D Thank you @dtan5457